# Daniel Iong, Matthew McAnear, Yuezhou Qu

### v1

@@ Acknowledgements
YC is supported by NSF DMS 2113397, NSF PHY 2027555, NASA 22-SWXC22_2-0005 and 22-SWXC22_2-0015.
@@ Appendix A Variational inference
Variational inference (VI) is commonly used to approximate posterior distributions in complex Bayesian models where the posterior is not tractable [38]. It provides an alternative to Markov chain Monte-Carlo (MCMC) sampling that tends to be more computationally efficient and scalable. VI approximates the posterior distribution by finding the distribution belonging to a simpler family of distributions that minimizes the Kullback-Leibler (KL) divergence with the exact posterior [38]. More concretely, suppose we have observed data, denoted by $\mathbf{y}\in\mathbb{R}^{n}$, and latent variables, denoted by $\mathbf{f}\in\mathbb{R}^{m}$. For simplicity, we assume that $\mathbf{y}$ and $\mathbf{f}$ are real-valued but that is not a requirement for VI. In the context of GPR, $\mathbf{f}$ are the latent function values. The goal of VI is to solve the following optimization problem:
@@
The resulting distribution $q^{*}(\mathbf{f})$ is then used as an approximation to the exact posterior. The difficulty of this optimization problem is dictated by the variational family $\mathcal{Q}$. Common choices include the mean-field and Gaussian variational family [39, 40]. The mean-field variational family consist of distributions that asssume the latent variables are mutually independent, i.e.
@@
The KL divergence depends on the marginal log likelihood $\log p(\mathbf{y})$ which is almost always intractable. We can rearrange eq. 28 to express the marginal log likelihood as
@@
This function is known as the evidence lower bound (ELBO) because it lower bounds the marginal log likelihood (or evidence). This follows from the KL divergence term being non-negative. Furthermore, since $\log p(\mathbf{y})$ is a constant with respect to $q$, maximizing the ELBO is equivalent to minimizing the KL divergence between the approximate and exact posterior, thereby solving the original optimization problem in eq. 26.
@@ Appendix C ANN Fitting
Due to computational constraints, hyperparameter tuning for the ANN model is conducted on the OTT station. Hyperparameter tuning station by station for the ANN model would require moving the neural net fitting code to a GPU, a path that was not explored for this paper. Therefore, we estimate that further enhancements to the ANN model’s performance are possible. In particular, the hyperparameter optimization performed for the OTT station may not generalize well to other stations, particularly those at higher latitudes. Nonetheless, the current model serves as an adequate baseline, demonstrating comparable performance with prior works such as [34] and [36] when comparing RMSE and skill scores across all test stations.
We opt for a 4-layer deep neural network architecture. Following common convention, we incorporate the rectified linear unit (ReLU) activation function and a batch size of 32. In multivariate regression problems with time series data, it is common to face the issue of gradient explosion. To mitigate this, we utilize the AdamW optimizer with a weight decay of 0.01 and implement gradient clipping with a maximum norm of 1.0. To further reduce overfitting, early stopping is employed after 5 epochs without improvement. The optimization of additional hyperparameters is facilitated by the Bayesian Optimization library [33], by randomly selecting 20 initial combinations of hyperparameters and conducting 40 further iterations. This procedure is iterated multiple times to identify a stable set of parameters. The learning rate is adjusted using an exponential decay schedule, and dropout layers are introduced following each hidden layer at the same rate. The weight of the loss function is also tuned in this process, so we utilize RMSE as the evaluation metric on our validation set. Hyperparameter tuning results are detailed in table 6.
@@
It is noteworthy that our ANN models tend to produce overly broad confidence intervals, particularly for extreme events and during storm periods. To address this issue, we predict a transformed scale parameter $\tilde{\sigma}^{2}$, with the final scale parameter $\hat{\sigma}^{2}$ calculated as
@@
where $A$, $B$, and $C$ are tunable hyperparameters ensuring that $\hat{\sigma}^{2}$ falls within the range $[0,C]$. After exploratory adjustments, these parameters are set to $A=0.01$, $B=500$, and $C=1000$. Though this change reduces interval length in extreme cases, it does not eliminate the problem of exploding interval lengths completely.

### v3

@@ Acknowledgements
This work was supported by the NSF SWQU project PHY 2027555. YC is further supported by NSF DMS 2113397, NASA 22-SWXC22_2-0005, and 22-SWXC22_2-0015.
@@ Appendix A Variational inference
Variational inference (VI) is commonly used to approximate posterior distributions in complex Bayesian models where the posterior is not tractable (Blei et al., 2017). It provides an alternative to Markov chain Monte-Carlo (MCMC) sampling that tends to be more computationally efficient and scalable. VI approximates the posterior distribution by finding the distribution belonging to a simpler family of distributions that minimizes the Kullback-Leibler (KL) divergence with the exact posterior (Blei et al., 2017). More concretely, suppose we have observed data, denoted by $\mathbf{y}\in\mathbb{R}^{n}$, and latent variables, denoted by $\mathbf{f}\in\mathbb{R}^{m}$. For simplicity, we assume that $\mathbf{y}$ and $\mathbf{f}$ are real-valued but that is not a requirement for VI. In the context of GPR, $\mathbf{f}$ are the latent function values. The goal of VI is to solve the following optimization problem:
@@
The resulting distribution $q^{*}(\mathbf{f})$ is then used as an approximation to the exact posterior. The difficulty of this optimization problem is dictated by the variational family $\mathcal{Q}$. Common choices include the mean-field and Gaussian variational family (Wainwright and Jordan, 2007; Opper and Archambeau, 2009). The mean-field variational family consists of distributions that asssumes the latent variables are mutually independent, i.e.
@@
The KL divergence depends on the marginal log-likelihood $\log p(\mathbf{y})$ which is almost always intractable. We can rearrange eq. 28 to express the marginal log-likelihood as
@@
This function is known as the evidence lower bound (ELBO) because it lower bounds the marginal log-likelihood (or evidence). This follows from the KL divergence term being non-negative. Furthermore, since $\log p(\mathbf{y})$ is a constant with respect to $q$, maximizing the ELBO is equivalent to minimizing the KL divergence between the approximate and exact posterior, thereby solving the original optimization problem in eq. 26.
@@ Appendix C ANN Fitting
Due to computational constraints, hyperparameter tuning for the ANN model is conducted on the OTT station. Hyperparameter tuning station by station for the ANN model would require moving the neural net fitting code to a GPU, a path that was not explored for this paper. Therefore, we estimate that further enhancements to the ANN model’s performance are possible. In particular, the hyperparameter optimization performed for the OTT station may not generalize well to other stations, particularly those at higher latitudes. Nonetheless, the current model serves as an adequate baseline, demonstrating comparable performance with prior works such as Keesee et al., (2020) and Pinto et al., (2022) when comparing RMSE and skill scores across all test stations.
We opt for a 4-layer deep neural network architecture. Following common convention, we incorporate the rectified linear unit (ReLU) activation function and a batch size of 32. In multivariate regression problems with time series data, it is common to face the issue of gradient explosion. To mitigate this, we utilize the AdamW optimizer with a weight decay of 0.01 and implement gradient clipping with a maximum norm of 1.0. To further reduce overfitting, early stopping is employed after 5 epochs without improvement. The following additional hyperparameters are tuned simultaneously by the Bayesian Optimization library Nogueira, 14 , by randomly selecting 20 initial combinations of hyperparameters and conducting 40 further iterations. This procedure is repeated multiple times to identify a stable set of hyperparameters. The learning rate is adjusted using an exponential decay schedule, and dropout layers are introduced following each hidden layer at the same rate. The weight of the loss function is also tuned in this process, so we utilize RMSE as the evaluation metric on our validation set. Hyperparameter tuning results are detailed in table 7.
@@
Our final model structure and performance results are similar to previous literature of the field (Keesee et al., 2020; Pinto et al., 2022). However, it is noteworthy that our ANN models tend to produce overly broad confidence intervals, particularly for extreme events and during storm periods. To address this issue, we predict a transformed scale parameter $\tilde{\sigma}^{2}$, with the final scale parameter $\hat{\sigma}^{2}$ calculated as
@@
where $A$, $B$, and $C$ are tunable hyperparameters ensuring that $\hat{\sigma}^{2}$ falls within the range $[0,C]$. After exploratory adjustments, these parameters are set to $A=0.01$, $B=500$, $C=1000$ for lower latitude stations (FRD, FRN, NEW, OTT, WNG), and $C=2000$ for higher latitude stations (ABK, HRN, IQA, MEA, PBQ, YKC). Though this change reduces interval length in extreme cases, it does not eliminate the problem of exploding interval lengths completely.
Table 8: Performance of an ANN model with MC-Dropout. All the parameters are unchanged, except for dropout rate, which is increased to 0.7 to improve its inference capability. 1000 samples are used to estimate the distribution of the final prediction.
Station RMSE Coverage Interval mean Interval median
NEW 16.786 0.742 23.435 16.043
OTT 16.006 0.747 24.291 16.551
WNG 14.398 0.748 22.332 15.693
YKC 96.641 0.621 86.609 65.661
ABK 89.685 0.618 72.297 43.356
MEA 78.873 0.646 41.405 24.306
FRN 14.138 0.640 20.136 14.555
IQA 82.633 0.583 75.083 61.914
PBQ 87.875 0.587 75.239 48.566
FRD 11.863 0.705 19.554 14.116
HRN 65.193 0.613 70.411 56.783
FUR 12.736 0.735 19.330 13.353
Table 9: Bias and variance of GPR-CN and ANN models for the testing period of 2015-2016. The ANN method predictions consistently have lower bias and lower variance, at the cost of prediction interval length.
Full Set Storm Only
Bias Variance Bias Variance
Model ANN GPR-CN ANN GPR-CN ANN GPR-CN ANN GPR-CN
Station
ABK 6.061 28.750 7685.095 6578.867 -2.257 43.218 21056.316 12067.962
FRD 2.830 4.467 132.663 150.201 12.588 16.637 473.219 726.362
FRN 3.851 7.558 181.500 184.933 8.518 23.736 1065.965 793.572
HRN 12.330 19.686 4094.969 3930.555 45.736 49.804 8905.957 8757.528
IQA -1.114 25.646 6605.617 6383.506 15.442 52.333 11859.860 13883.121
MEA 10.806 20.898 5275.414 5763.874 21.143 42.103 12670.594 15761.465
NEW 3.277 6.254 258.872 309.885 10.275 15.154 1012.890 1494.153
OTT 2.492 4.901 234.728 317.216 10.450 17.275 1553.697 2806.177
PBQ 12.647 34.225 7184.451 7336.389 25.398 49.459 12220.211 14443.320
WNG 2.288 4.715 188.606 202.985 9.351 15.978 787.910 971.462
YKC 5.450 26.849 8761.938 8617.623 16.761 45.561 20696.077 23118.061